OpenStudy (anonymous):

if you have endpoints of a diameter (0,0) and (6,8) and you have to write the standard form of the equation of a circle, what would the origin be? i did the question assuming the origin is (0,0) and got the equation x^2 +y^2=25. but apparently thats wrong and the equation is (x-3)^2 + (y-4)^2 =25.. meaning that the origin is assumed to be (3,4).... basically how would one know what the origin is in this instance as both equations are diff?

OpenStudy (anonymous):

by "origin" do you mean "center"?

OpenStudy (anonymous):

if (0,0) is one endpoint of the diameter it is not also the center. since the other endpoint is (6,8) the center is half way between i.e. at the point (3,4)

OpenStudy (anonymous):

yes i mean center of the circle

OpenStudy (anonymous):

omg okay thanks blond moment :)

OpenStudy (anonymous):

so your equation will look like \[(x-3)^2+(y-4)^2=r^2\] and you find \(r^2\) by pythagoras

OpenStudy (anonymous):

square of the distance between (0,0) and (3,4) is \(3^2+4^2=25\) gentlemen prefer blondes

OpenStudy (anonymous):

haha yeah yeah. thanks for the help :)